# Uniform approximation on certain polynomial polyhedra in $\mathbb{C}^{2}$

### v1

@@ 2. Technical Results
[…]
@@
Assume that there is a subsequence $r_{j_{k}}$ of $r_{j}$ such that for each $k$ there exists a point ${\sf Gr}_{h}(a_{k},b_{k})\in\widehat{{\sf Gr}_{h}(\mathfrak{Q}_{j_{k}})}$ with $|p_{1}(a_{k},b_{k})|=r_{j_{k}}$ and $|p_{2}(a_{k},b_{k})|<r_{j_{k}}-\varepsilon_{j_{k}}.$ Then there exists a analytic set normal-$\triangle$ and all $h_{j}$’s are holomorphic on .normal-$\triangle$.△ .
@@
To see a proof of the claim, let $\sup_{{\sf Gr}_{h}(\mathfrak{Q}_{j})}|\kappa_{k}(z,w)|=|\kappa_{k}(s,h(s))|,$ for some $s=(s_{1},s_{2})\in\mathfrak{Q}_{j}.$ If $s\in\triangle_{k},$ then from (3.7) and (3.8), we get that $|\kappa_{k}(s,h(s))|<1.$ If $s\in\mathfrak{Q}_{j}\setminus\triangle_{k},$ then for large enough $m,$ we have that $|\kappa_{k}(s,h(s))|<1.$ This is a contradiction. Hence ${\sf Gr}_{h}(a_{k},b_{k})\in\widehat{{\sf Gr}_{h}(\triangle_{k})}.$ This implies ${\sf Gr}_{u}(a_{k},b_{k})\in\widehat{{\sf Gr}_{u}(\triangle_{k})},$ where $u$ is given by 3.5. Note that $\{p_{1}(z)=e^{i\theta_{0}}p_{1}(a_{k},b_{k})\}$ is equivalent to $\mathbb{C}$ (by Result 2.8), and $u_{j}$’s are harmonic on $\triangle_{k}.$ Hence $u$ is constant on some component of $\triangle_{k},$ i.e. all $h_{j}$’s are holomorphic on that component of $\triangle_{k}.$ We will denote this component again by $\triangle_{k}.$ We now claim that ${\sf Gr}_{h}(\triangle_{k})\subset\widehat{{\sf Gr}_{h}(\{p_{1}(z)=e^{i\theta_{0}}p_{1}(a_{k},b_{k}),|p_{2}(z)|=r_{j_{k}}\})}.$ To prove this claim, let $(\xi,h(\xi))\in{\sf Gr}_{h}(\triangle_{k})$ for some $\xi\in\triangle_{k}.$ Then for any polynomial $P(z,w)$ in $\mathbb{C}^{2},$
@@
Since ${\sf Gr}_{u}(\alpha_{0})\in\widehat{{\sf Gr}_{u}(\mathfrak{Q}_{j})}$ and ${\sf Gr}_{u}(\alpha_{0})\in F^{-1}\{0\},$ hence ${\sf Gr}_{u}(\alpha_{0})\in\widehat{{\sf Gr}_{u}(\mathfrak{Q}^{{\bf c}}_{j})}.$ Now,
@@
If $\alpha_{0}\in L_{c_{1}}\setminus W,$ we are done. If $\widehat{\mathfrak{Q}^{\bf c}_{j}}$ contains only singular points and since singular points are isolated, $\widehat{\mathfrak{Q}^{\bf c}_{j}}$ is finite. Therefore, $\alpha_{0}\in\widehat{\mathfrak{Q}^{\bf c}_{j}}=\mathfrak{Q}^{\bf c}_{j},$ but we have taken $\alpha_{0}$ from $\mathfrak{D}^{r_{j}}_{2}.$ This is a contradiction. Therefore, $\widehat{\mathfrak{Q}^{\bf c}_{j}}$ contains at least one point $z_{0}\in L_{c_{1}}\setminus W.$ We claim that ${\sf Gr}_{u}(z_{0})\in\widehat{{\sf Gr}_{u}(\mathfrak{Q}_{j})}.$ To prove this claim, let $H(z,w)$ be a polynomial in $\mathbb{C}^{2+N}.$ Since $(z_{0},u(z_{0}))\in{\sf Gr}_{u}(\widehat{\mathfrak{Q}^{\bf c}_{j}})$, then
@@
Case I If $s\in\mathfrak{L}_{z_{0}}\cap\partial U$ and since $s\in X,$ therefore $s\in X_{z_{0}}=X\cap\partial D(z_{0}),$ then we have $|\gamma(s)|<1.$ Also, since $\Phi_{1}(s)\in i\mathbb{R},$ there exists $y\in\mathbb{R}$ such that $\Phi_{1}(s)=iy.$ Therefore, $(\xi\circ\Phi_{1}(s))=e^{-y^{2}}<1.$ This implies $|(\xi\circ\Phi_{1}(s))^{d}|\leq 1.$ This contradicts that $z_{0}\in\widehat{X}.$ Case II If $s\in(L_{\bf c}\cap\bar{U}\setminus D(z_{0}))$ and $s\notin\mathfrak{L}_{z_{0}}\cap\partial U$ implies $s\notin\mathfrak{L}_{z_{0}},$ by (3.10), $|\xi\circ\Phi_{1}(s))|<1.$ Then for sufficiently large $d$, we have $|\gamma(s)|\times|\xi\circ\Phi_{1}(s))|^{d}<1.$ Again this is a contradiction to the fact that $z_{0}\in\widehat{X}.$ Therefore, $X_{z_{0}}=\partial D(z_{0})\subset X.$ Hence,
@@
Since, by assumption in the beginning of the proof, ${\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})$ is not polynomially convex, by (3.4), there exist $\alpha_{0}\in\widetilde{V}$ such that ${\sf Gr}_{h}(\alpha_{0})\in\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}.$ Case I: Assume that $\widetilde{V}\not=\mathfrak{D}_{2}.$ Then Lemma 3.2 gives an irreducible component $V$ of $\widetilde{V}$ such that all $h_{j}$’s are holomorphic on $V.$ Observe that $\overline{V}\setminus V\subset\Gamma_{\mathfrak{D}_{2}}.$ The proof of the theorem is complete in this case in view of Lemma 2.22. Case II: Now we consider the case that $\widetilde{V}=\mathfrak{D}_{2}.$ Without loss of generality, we assume that $h_{1}$ is non-holomorphic. Since ${\sf Gr}_{h}(\alpha_{0})\in\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})},$ there exists $j_{0}\in\mathbb{N}$ such that $\alpha_{0}\in\mathfrak{D}^{r_{j}}_{2}$ for $j>j_{0}.$ Therefore, by (3.6) and Result 2.4, we obtain that ${\sf Gr}_{h}(\alpha_{0})\in\widehat{{\sf Gr}_{h}(\mathfrak{Q}_{j})}:$ in fact, assume that
@@ 4. Description of hull
If #$\#CC(X)=\infty,$ then there is nothing to prove. In view Result 2.11, it is not possible that #$\#CC(X)<\infty,$ but #$\#CC(\widehat{X})=\infty.$ In fact, assume that $\widehat{X}=\cup^{\infty}_{j=1}X_{j},$ where each $X_{j}$ is disjoint connected component of $\widehat{X}.$ This implies $X=X\cap\widehat{X}=\cup^{\infty}_{j=1}X_{j}\cap X.$ Result 2.11 says that $X_{j}=\widehat{X_{j}\cap X}~{}\forall j\geq 1.$ Since each $X_{j}\cap X$ is disjoint, and #$\#CC(X)<\infty,$ therefore there exists $j_{0}\in\mathbb{N}$ such that $X_{j}\cap X=\emptyset~{}\forall j>j_{0}.$ Hence $X_{j}=\widehat{X_{j}\cap X}=\emptyset~{}\forall j>j_{0}.$ This is a contradiction to the fact that #$\#CC(\widehat{X})=\infty.$ Next, we assume that #$\#CC(X)=m$ and #$\#CC(\widehat{X})=l.$ Assume that $\widehat{X}=\cup^{l}_{j=1}X_{j},$ where each $X_{j}$ is disjoint connected component of $\widehat{X}.$ Then $X=\widehat{X}\cap X=\cup^{l}_{j=1}X_{j}\cap X.$ By using Result 2.11, we obtain that $X_{j}=\widehat{X_{j}\cap X}~{}\forall j\geq 1.$ If each $X_{j}\cap X$ is connected, then $l=m.$ Otherwise $l<m.$
@@
for some $l\in\mathbb{N},$ where each $q_{j}$ is an irreducible polynomial in $\mathbb{C}^{2}.$ We define the corresponding irreducible algebraic variety $Z_{j}:=\{z\in\mathbb{C}^{2}:q_{j}(z)=0\}.$ We assume $\triangle({z})\not\equiv 0$ on $X.$ Therefore, each $q_{j}$ is a non-zero holomorphic polynomial in $\mathbb{C}^{2}.$ We set
@@
Assume that $\sup_{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{{\mathfrak{D}_{2}}})}|g(z_{1},z_{2},\xi)|=|g(s,\overline{P\circ\Psi}(s))|$ for some $s=(s_{1},s_{2})\in\Gamma_{{\mathfrak{D}_{2}}}.$ Case I: $s\in\Gamma_{{\mathfrak{D}_{2}}}\setminus(U\cup W),$ then there is nothing to prove. Case II: $s\in U$ and $s\in\Gamma_{{\mathfrak{D}_{2}}}.$ Then $|g_{1}(s,\overline{P\circ\Psi}(s))|\leq 1$ but $|p_{0}(s)f_{2}(s)|<\frac{1}{2}.$ This implies $|g(s,\overline{P\circ\Psi}(s))|<\frac{1}{2}.$ Case III: $s\in W$ and $s\in\Gamma_{{\mathfrak{D}_{2}}}.$ Then $|g_{1}(s,\overline{P\circ\Psi}(s))|\leq 1$ but $|p_{0}(s)f_{2}(s)|<\frac{1}{2}.$ This implies $|g(s,\overline{P\circ\Psi}(s))|<\frac{1}{2}.$ Combining all the three cases, we can conclude that ${\sf Gr}_{\overline{(P\circ\Psi)}}(a_{0})\notin\widehat{{\sf Gr}_{\overline{(P\circ\Psi)}}(\Gamma_{{\mathfrak{D}_{2}}})}.$ Next, we assume that $a_{0}\in(Z_{j_{0}}\cap\overline{\mathfrak{D}}_{2}\cap A)\setminus(L\cup\Gamma_{{\mathfrak{D}_{2}}}).$ Since $(Z_{j_{0}}\cap\overline{\mathfrak{D}}_{2}\cap A)\setminus(L\cup\Gamma_{{\mathfrak{D}_{2}}})$ is finite and $\widehat{{\sf Gr}_{\overline{(P\circ\Psi)}}(\Gamma_{{\mathfrak{D}_{2}}})}$ does not containing any isolated point, therefore, in this case also ${\sf Gr}_{\overline{(P\circ\Psi)}}(a_{0})\notin\widehat{{\sf Gr}_{\overline{(P\circ\Psi)}}(\Gamma_{{\mathfrak{D}_{2}}})}.$
@@ 5. Examples
or there exists a non-trivial algebraic variety $V\subset\mathbb{C}^{2}$ with $V\cap\partial\mathfrak{D}_{2}\subset\Gamma_{\mathfrak{D}_{2}},$ and the pluriharmonic extensions of the $f_{j}$ are holomorphic on $V.$ Explanation: We claim that
@@
Proof of (iv): Without loss of generality, assume that $|a|\neq 0\neq|c|.$ Then $p_{1}(z)=e_{1}$ and $p_{2}(z)=e_{2}$ are simply connected for any $e_{1},e_{2}\in\mathbb{C}$. To prove this we use the following change of coordinate system $(z^{*}_{1},z^{*}_{2})=(az_{1}+bz^{l}_{2}-e_{1},z_{2})$ for the first one and $(z^{*}_{1},z^{*}_{2})=(cz_{1}+dz^{l}_{2}-e_{2},z_{2})$ for the second one. Then $p_{1}(z)=e_{1}$ and $p_{2}(z)=e_{2}$ becomes $\{(z^{*}_{1},z^{*}_{2})\in\mathbb{C}^{2}:z^{*}_{1}=0\}$ and $\{(z^{*}_{1},z^{*}_{2})\in\mathbb{C}^{2}:z^{*}_{1}=0\}$ respectively. Now by using Theorem 1.2, we can conclude that either
@@
Explanation: Following the notation in Theorem 4.3, let $\Psi(z)=(p_{1}(z),p_{2}(z)),$ $h(z)=\frac{1}{z_{1}}+\frac{1}{z_{2}}.$ Therefore,
@@
We define $q_{1}:=p_{1}-p_{2}=z_{1}+z_{2}-z_{1}z_{2},~{}~{}q_{2}:=p_{1}+p_{2}=z_{1}+z_{2}+z_{1}z_{2},~{}~{}q_{3}=z_{2}-z_{1}$ and $Z_{j}=\{q_{j}=0\}.$ Therefore, Clearly, $Z_{j}\cap\overline{\mathfrak{D}}_{2}\setminus(\Gamma_{{\mathfrak{D}_{2}}}\cup L)\subset X$ only for ${j}=2.$ It is easy to see that $\widehat{Q}_{2}=\widehat{Z_{2}\cap\Gamma_{{\mathfrak{D}_{2}}}}=\{(z_{1},z_{2})\in\overline{\mathfrak{D}}_{2}:z_{1}+z_{2}+z_{1}z_{2}=0\}.$ Indeed, we write $\mathcal{K}=\{(z_{1},z_{2})\in\mathbb{C}^{2}:z_{1}+z_{2}=0,|z_{1}|=1,|z_{2}|=1\}.$ Clearly, $\Psi^{-1}(\mathcal{K})=Q_{2}=Z_{2}\cap\Gamma_{{\mathfrak{D}_{2}}}.$ We now claim that
@@
Hence $\widehat{\mathcal{K}}=\widehat{\Psi(Q_{2})}=\widehat{\Psi(\widehat{Q_{2}})}.$ Since $\Psi^{-1}(\Psi(\widehat{Q_{2}}))=\widehat{Q_{2}}$ is polynomially convex and $\Psi$ is proper map, therefore $\Psi(\widehat{Q_{2}})$ is also polynomially convex. Thus $\widehat{\Psi(Q_{2})}=\widehat{\Psi(\widehat{Q_{2}})}=\Psi(\widehat{Q_{2}}).$ This implies $\widehat{Q_{2}}=\Psi^{-1}(\widehat{\mathcal{K}}),$ i.e. $\widehat{Q_{2}}=\overline{\mathfrak{D}}_{2}\cap Z_{2}.$ Therefore, by Theorem 4.3, we get that
@@
We define $q_{1}:=p_{1}-p_{2}=z_{1}+2z^{2}_{2},~{}~{}q_{2}:=p_{1}+p_{2}=3z_{1},~{}~{}q_{3}=-6z_{2}$ and $Z_{j}=\{q_{j}=0\}.$ Clearly, $Q_{3}=Z_{3}\cap\Gamma_{{\mathfrak{D}_{2}}}=\emptyset$ and it is easy to show that if $z\in Z_{2}\cap X,$ then $z\in\Gamma_{{\mathfrak{D}_{2}}},$ this is not possible because $X\cap\Gamma_{{\mathfrak{D}_{2}}}=\emptyset.$ Now its remaining to verify for the set $Q_{1}=Z_{1}\cap\Gamma_{{\mathfrak{D}_{2}}}.$ It is easy to see that $Z_{1}\cap\overline{\mathfrak{D}}_{2}\setminus(\Gamma_{{\mathfrak{D}_{2}}}\cup L)\subset X$ and $\widehat{Q}_{1}=\widehat{Z_{1}\cap\Gamma_{{\mathfrak{D}_{2}}}}=\{(z_{1},z_{2})\in\overline{\mathfrak{D}}_{2}:z_{1}+2z^{2}_{2}=0\}.$ Therefore, by Theorem 4.3, we get that
@@
Therefore, by using Theorem 2.24, we get $[z_{1},z_{2},\overline{(P\circ\Psi)}(z);\Gamma_{\mathfrak{D}_{2}}]=\mathcal{C}(\Gamma_{\mathfrak{D}_{2}}).$ Acknowledgments. The first named author was partially supported by a MATRICS Research Grant (MTR/2017/000974) of SERB, Dept. of Science and Technology, Govt. of India, for the beginning of this work and is supported partially by a Core Research Grant (CRG/2022/003560) of SERB, Dept. of Science and Technology, Govt. of India, for the later part of the work. The second named author acknowledges partial support from an INSPIRE Fellowship (IF 160487) awarded by the DST, Govt. of India, for the early version of this work, and an IoE-IISc Postdoctoral Fellowship (R(HR)(IOE-PDF)(MA)(GMM)-114) for the latest version of this work. The second author would also like to thank Amar Deep Sarkar for the fruitful discussions.

### v2

@@ 2. Technical Results
Assume that $\widetilde{V}\not=\mathfrak{D}_{2}.$ Let $V$ be an irreducible component of $\widetilde{V}$ and let $\alpha\in V\cap(\widetilde{V})_{\text{reg}}$ with ${\sf Gr}_{h}(\alpha)\in\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}.$ Then ${\sf Gr}_{h}(V)\subset\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}$ and all $h_{j}$’s are holomorphic on $V.$
@@
Assume that there is a subsequence $r_{j_{k}}$ of $r_{j}$ such that for each $k$ there exists a point ${\sf Gr}_{h}(a_{k},b_{k})\in\widehat{{\sf Gr}_{h}(\mathfrak{Q}_{j_{k}})}$ with $|p_{1}(a_{k},b_{k})|=r_{j_{k}}$ and $|p_{2}(a_{k},b_{k})|<r_{j_{k}}-\varepsilon_{j_{k}}.$ Then there exists a analytic set $\triangle$ and all $h_{j}$’s are holomorphic on $\triangle$.△ .
@@
To see a proof of the claim, let $\sup_{{\sf Gr}_{h}(\mathfrak{Q}_{j})}|\kappa_{k}(z,w)|=|\kappa_{k}(s,h(s))|,$ for some $s=(s_{1},s_{2})\in\mathfrak{Q}_{j}.$ If $s\in\triangle_{k},$ then from (3.7) and (3.8), we get that $|\kappa_{k}(s,h(s))|<1.$ If $s\in\mathfrak{Q}_{j}\setminus\triangle_{k},$ then for large enough $m,$ we have that $|\kappa_{k}(s,h(s))|<1.$ This is a contradiction. Hence ${\sf Gr}_{h}(a_{k},b_{k})\in\widehat{{\sf Gr}_{h}(\triangle_{k})}.$ This implies ${\sf Gr}_{u}(a_{k},b_{k})\in\widehat{{\sf Gr}_{u}(\triangle_{k})},$ where $u$ is given by 3.5. Note that $\{p_{1}(z)=e^{i\theta_{0}}p_{1}(a_{k},b_{k})\}$ is equivalent to $\mathbb{C}$ (by Result 2.8), and $u_{j}$’s are harmonic on $\triangle_{k}.$ Hence $u$ is constant on some component of $\triangle_{k},$ i.e. all $h_{j}$’s are holomorphic on that component of $\triangle_{k}.$ We will denote this component again by $\triangle_{k}.$
We now claim that ${\sf Gr}_{h}(\triangle_{k})\subset\widehat{{\sf Gr}_{h}(\{p_{1}(z)=e^{i\theta_{0}}p_{1}(a_{k},b_{k}),|p_{2}(z)|=r_{j_{k}}\})}.$ To prove this claim, let $(\xi,h(\xi))\in{\sf Gr}_{h}(\triangle_{k})$ for some $\xi\in\triangle_{k}.$ Then for any polynomial $P(z,w)$ in $\mathbb{C}^{2},$
@@
Since ${\sf Gr}_{u}(\alpha_{0})\in\widehat{{\sf Gr}_{u}(\mathfrak{Q}_{j})}$ and ${\sf Gr}_{u}(\alpha_{0})\in F^{-1}\{0\},$ hence ${\sf Gr}_{u}(\alpha_{0})\in\widehat{{\sf Gr}_{u}(\mathfrak{Q}^{{\bf c}}_{j})}.$
Now,
@@
If $\alpha_{0}\in L_{c_{1}}\setminus W,$ we are done. If $\widehat{\mathfrak{Q}^{\bf c}_{j}}$ contains only singular points and since singular points are isolated, $\widehat{\mathfrak{Q}^{\bf c}_{j}}$ is finite. Therefore, $\alpha_{0}\in\widehat{\mathfrak{Q}^{\bf c}_{j}}=\mathfrak{Q}^{\bf c}_{j},$ but we have taken $\alpha_{0}$ from $\mathfrak{D}^{r_{j}}_{2}.$ This is a contradiction. Therefore, $\widehat{\mathfrak{Q}^{\bf c}_{j}}$ contains at least one point $z_{0}\in L_{c_{1}}\setminus W.$
We claim that ${\sf Gr}_{u}(z_{0})\in\widehat{{\sf Gr}_{u}(\mathfrak{Q}_{j})}.$ To prove this claim, let $H(z,w)$ be a polynomial in $\mathbb{C}^{2+N}.$ Since $(z_{0},u(z_{0}))\in{\sf Gr}_{u}(\widehat{\mathfrak{Q}^{\bf c}_{j}})$, then
@@
Case I If $s\in\mathfrak{L}_{z_{0}}\cap\partial U$ and since $s\in X,$ therefore $s\in X_{z_{0}}=X\cap\partial D(z_{0}),$ then we have $|\gamma(s)|<1.$ Also, since $\Phi_{1}(s)\in i\mathbb{R},$ there exists $y\in\mathbb{R}$ such that $\Phi_{1}(s)=iy.$ Therefore, $(\xi\circ\Phi_{1}(s))=e^{-y^{2}}<1.$ This implies $|(\xi\circ\Phi_{1}(s))^{d}|\leq 1.$ This contradicts that $z_{0}\in\widehat{X}.$
Case II If $s\in(L_{\bf c}\cap\bar{U}\setminus D(z_{0}))$ and $s\notin\mathfrak{L}_{z_{0}}\cap\partial U$ implies $s\notin\mathfrak{L}_{z_{0}},$ by (3.10), $|\xi\circ\Phi_{1}(s))|<1.$ Then for sufficiently large $d$, we have $|\gamma(s)|\times|\xi\circ\Phi_{1}(s))|^{d}<1.$ Again this is a contradiction to the fact that $z_{0}\in\widehat{X}.$ Therefore, $X_{z_{0}}=\partial D(z_{0})\subset X.$ Hence,
@@
Since, by assumption in the beginning of the proof, ${\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})$ is not polynomially convex, by (3.4), there exist $\alpha_{0}\in\widetilde{V}$ such that ${\sf Gr}_{h}(\alpha_{0})\in\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}.$
Case I: Assume that $\widetilde{V}\not=\mathfrak{D}_{2}.$ Then Lemma 3.2 gives an irreducible component $V$ of $\widetilde{V}$ such that all $h_{j}$’s are holomorphic on $V.$ Observe that $\overline{V}\setminus V\subset\Gamma_{\mathfrak{D}_{2}}.$ The proof of the theorem is complete in this case in view of Lemma 2.22.
Case II: Now we consider the case that $\widetilde{V}=\mathfrak{D}_{2}.$ Without loss of generality, we assume that $h_{1}$ is non-holomorphic. Since ${\sf Gr}_{h}(\alpha_{0})\in\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})},$ there exists $j_{0}\in\mathbb{N}$ such that $\alpha_{0}\in\mathfrak{D}^{r_{j}}_{2}$ for $j>j_{0}.$ Therefore, by (3.6) and Result 2.4, we obtain that ${\sf Gr}_{h}(\alpha_{0})\in\widehat{{\sf Gr}_{h}(\mathfrak{Q}_{j})}:$ in fact, assume that
@@ 4. Description of hull
If #$\#CC(X)=\infty,$ then there is nothing to prove. In view Result 2.11, it is not possible that #$\#CC(X)<\infty,$ but #$\#CC(\widehat{X})=\infty.$ In fact, assume that $\widehat{X}=\cup^{\infty}_{j=1}X_{j},$ where each $X_{j}$ is disjoint connected component of $\widehat{X}.$ This implies $X=X\cap\widehat{X}=\cup^{\infty}_{j=1}X_{j}\cap X.$ Result 2.11 says that $X_{j}=\widehat{X_{j}\cap X}~{}\forall j\geq 1.$ Since each $X_{j}\cap X$ is disjoint, and #$\#CC(X)<\infty,$ therefore there exists $j_{0}\in\mathbb{N}$ such that $X_{j}\cap X=\emptyset~{}\forall j>j_{0}.$ Hence $X_{j}=\widehat{X_{j}\cap X}=\emptyset~{}\forall j>j_{0}.$ This is a contradiction to the fact that #$\#CC(\widehat{X})=\infty.$
Next, we assume that #$\#CC(X)=m$ and #$\#CC(\widehat{X})=l.$ Assume that $\widehat{X}=\cup^{l}_{j=1}X_{j},$ where each $X_{j}$ is disjoint connected component of $\widehat{X}.$ Then $X=\widehat{X}\cap X=\cup^{l}_{j=1}X_{j}\cap X.$ By using Result 2.11, we obtain that $X_{j}=\widehat{X_{j}\cap X}~{}\forall j\geq 1.$ If each $X_{j}\cap X$ is connected, then $l=m.$ Otherwise $l<m.$
@@
for some $l\in\mathbb{N},$ where each $q_{j}$ is an irreducible polynomial in $\mathbb{C}^{2}.$ We define the corresponding irreducible algebraic variety $Z_{j}:=\{z\in\mathbb{C}^{2}:q_{j}(z)=0\}.$ We assume $\triangle({z})\not\equiv 0$ on $X.$ Therefore, each $q_{j}$ is a non-zero holomorphic polynomial in $\mathbb{C}^{2}.$
We set
@@
Assume that $\sup_{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{{\mathfrak{D}_{2}}})}|g(z_{1},z_{2},\xi)|=|g(s,\overline{P\circ\Psi}(s))|$ for some $s=(s_{1},s_{2})\in\Gamma_{{\mathfrak{D}_{2}}}.$
Case I: $s\in\Gamma_{{\mathfrak{D}_{2}}}\setminus(U\cup W),$ then there is nothing to prove.
Case II: $s\in U$ and $s\in\Gamma_{{\mathfrak{D}_{2}}}.$ Then $|g_{1}(s,\overline{P\circ\Psi}(s))|\leq 1$ but $|p_{0}(s)f_{2}(s)|<\frac{1}{2}.$ This implies $|g(s,\overline{P\circ\Psi}(s))|<\frac{1}{2}.$
Case III: $s\in W$ and $s\in\Gamma_{{\mathfrak{D}_{2}}}.$ Then $|g_{1}(s,\overline{P\circ\Psi}(s))|\leq 1$ but $|p_{0}(s)f_{2}(s)|<\frac{1}{2}.$ This implies $|g(s,\overline{P\circ\Psi}(s))|<\frac{1}{2}.$
Combining all the three cases, we can conclude that ${\sf Gr}_{\overline{(P\circ\Psi)}}(a_{0})\notin\widehat{{\sf Gr}_{\overline{(P\circ\Psi)}}(\Gamma_{{\mathfrak{D}_{2}}})}.$
Next, we assume that $a_{0}\in(Z_{j_{0}}\cap\overline{\mathfrak{D}}_{2}\cap A)\setminus(L\cup\Gamma_{{\mathfrak{D}_{2}}}).$ Since $(Z_{j_{0}}\cap\overline{\mathfrak{D}}_{2}\cap A)\setminus(L\cup\Gamma_{{\mathfrak{D}_{2}}})$ is finite and $\widehat{{\sf Gr}_{\overline{(P\circ\Psi)}}(\Gamma_{{\mathfrak{D}_{2}}})}$ does not containing any isolated point, therefore, in this case also ${\sf Gr}_{\overline{(P\circ\Psi)}}(a_{0})\notin\widehat{{\sf Gr}_{\overline{(P\circ\Psi)}}(\Gamma_{{\mathfrak{D}_{2}}})}.$
@@ 5. Examples
Explanation: We claim that
@@
Proof of (iii): It is easy to see that $\overline{\mathfrak{D}}_{2}$ is path-connected. To see this, let $(a,b)\in\overline{\mathfrak{D}}_{2}.$ Then $\eta(t):=(t^{l}a,tb):[0,1]\to\overline{\mathfrak{D}}_{2}$ is a path between $(a,b)$ and the origin. Therefore, there is always a path between the origin and the arbitrary point of $\overline{\mathfrak{D}}_{2}.$ Hence $\overline{\mathfrak{D}}_{2}$ is path connected.
Now we show that $\overline{\mathfrak{D}}_{2}$ is simply-connected. For this, let $\gamma=(\gamma_{1},\gamma_{2}):[0,1]\to\overline{\mathfrak{D}}_{2}$ be a loop in $\overline{\mathfrak{D}}_{2}$ based at $(0,0).$ We consider the function $H(t,x):[0,1]\times[0,1]\to\overline{\mathfrak{D}}_{2}$ by
@@
Proof of (iv): Without loss of generality, assume that $|a|\neq 0\neq|c|.$ Then $p_{1}(z)=e_{1}$ and $p_{2}(z)=e_{2}$ are simply connected for any $e_{1},e_{2}\in\mathbb{C}$. To prove this we use the following change of coordinate system $(z^{*}_{1},z^{*}_{2})=(az_{1}+bz^{l}_{2}-e_{1},z_{2})$ for the first one and $(z^{*}_{1},z^{*}_{2})=(cz_{1}+dz^{l}_{2}-e_{2},z_{2})$ for the second one. Then $p_{1}(z)=e_{1}$ and $p_{2}(z)=e_{2}$ becomes $\{(z^{*}_{1},z^{*}_{2})\in\mathbb{C}^{2}:z^{*}_{1}=0\}$ and $\{(z^{*}_{1},z^{*}_{2})\in\mathbb{C}^{2}:z^{*}_{1}=0\}$ respectively.
Now by using Theorem 1.2, we can conclude that either
@@
Explanation: Following the notations in Theorem 4.3, let $\Psi(z)=(p_{1}(z),p_{2}(z)),$ $h(z)=\frac{1}{z_{1}}+\frac{1}{z_{2}}.$ Therefore,
@@
We define $q_{1}:=p_{1}-p_{2}=z_{1}+z_{2}-z_{1}z_{2},~{}~{}q_{2}:=p_{1}+p_{2}=z_{1}+z_{2}+z_{1}z_{2},~{}~{}q_{3}=z_{2}-z_{1}$ and $Z_{j}=\{q_{j}=0\}.$ Therefore,
Clearly, $Z_{j}\cap\overline{\mathfrak{D}}_{2}\setminus(\Gamma_{{\mathfrak{D}_{2}}}\cup L)\subset X$ only for ${j}=2.$ It is easy to see that $\widehat{Q}_{2}=\widehat{Z_{2}\cap\Gamma_{{\mathfrak{D}_{2}}}}=\{(z_{1},z_{2})\in\overline{\mathfrak{D}}_{2}:z_{1}+z_{2}+z_{1}z_{2}=0\}.$ Indeed, we write $\mathcal{K}=\{(z_{1},z_{2})\in\mathbb{C}^{2}:z_{1}+z_{2}=0,|z_{1}|=1,|z_{2}|=1\}.$ Clearly, $\Psi^{-1}(\mathcal{K})=Q_{2}=Z_{2}\cap\Gamma_{{\mathfrak{D}_{2}}}.$ We now claim that
@@
Hence $\widehat{\mathcal{K}}=\widehat{\Psi(Q_{2})}=\widehat{\Psi(\widehat{Q_{2}})}.$ Since $\Psi^{-1}(\Psi(\widehat{Q_{2}}))=\widehat{Q_{2}}$ is polynomially convex and $\Psi$ is proper map, therefore $\Psi(\widehat{Q_{2}})$ is also polynomially convex. Thus $\widehat{\Psi(Q_{2})}=\widehat{\Psi(\widehat{Q_{2}})}=\Psi(\widehat{Q_{2}}).$ This implies $\widehat{Q_{2}}=\Psi^{-1}(\widehat{\mathcal{K}}),$ i.e. $\widehat{Q_{2}}=\overline{\mathfrak{D}}_{2}\cap Z_{2}.$
Therefore, by Theorem 4.3, we get that
@@
We define $q_{1}:=p_{1}-p_{2}=z_{1}+2z^{2}_{2},~{}~{}q_{2}:=p_{1}+p_{2}=3z_{1},~{}~{}q_{3}=-6z_{2}$ and $Z_{j}=\{q_{j}=0\}.$
Clearly, $Q_{3}=Z_{3}\cap\Gamma_{{\mathfrak{D}_{2}}}=\emptyset$ and it is easy to show that if $z\in Z_{2}\cap X,$ then $z\in\Gamma_{{\mathfrak{D}_{2}}},$ this is not possible because $X\cap\Gamma_{{\mathfrak{D}_{2}}}=\emptyset.$ Now its remaining to verify for the set $Q_{1}=Z_{1}\cap\Gamma_{{\mathfrak{D}_{2}}}.$ It is easy to see that $Z_{1}\cap\overline{\mathfrak{D}}_{2}\setminus(\Gamma_{{\mathfrak{D}_{2}}}\cup L)\subset X$ and $\widehat{Q}_{1}=\widehat{Z_{1}\cap\Gamma_{{\mathfrak{D}_{2}}}}=\{(z_{1},z_{2})\in\overline{\mathfrak{D}}_{2}:z_{1}+2z^{2}_{2}=0\}.$ Therefore, by Theorem 4.3, we get that
@@
Explanation: Following the notations in Theorem 4.3, let $\Psi(z)=(p_{1}(z),p_{2}(z)),$ $h(z)=\frac{1}{z_{1}}+\frac{2}{z_{2}}.$ Therefore,
@@
Therefore, by using Theorem 2.24, we get $[z_{1},z_{2},\overline{(P\circ\Psi)}(z);\Gamma_{\mathfrak{D}_{2}}]=\mathcal{C}(\Gamma_{\mathfrak{D}_{2}}).$
Acknowledgments. The first named author was partially supported by a MATRICS Research Grant (MTR/2017/000974) of SERB, Dept. of Science and Technology, Govt. of India, for the beginning of this work and is supported partially by a Core Research Grant (CRG/2022/003560) of ANRF (formerly SERB), Govt. of India, for the later part of the work. The second named author acknowledges partial support from an INSPIRE Fellowship (IF 160487) awarded by the DST, Govt. of India, for the early version of this work, and an IoE-IISc Postdoctoral Fellowship (R(HR)(IOE-PDF)(MA)(GMM)-114) for the latest version of this work. The second named author would also like to thank Amar Deep Sarkar for the fruitful discussions.